Solve the following equation. Separate your answers with commas. Repeated roots should only be entered once.
1/x-7 + 1/x-2 = 9/14

Answer by stanbon(75887)

Solve the following equation. Separate your answers with commas. Repeated roots should only be entered once.
1/x-7 + 1/x-2 = 9/14
------------------------
Multiply thru by 14(x-2)(x-7) to get:
14(x-2) + 14(x-7) = (x-2)(x-7)
------
28x -126 = x^2 - 9x + 14
------
x^2 - 37x + 140 = 0
-------------
x = 32.72... or x = 4.28...
=============
Cheers,
Stan H.
-------------